The government of a poor developing country fears that a political upheaval will occur unless



the growth rate is at least 6 percent per annum. The ICOR and the saving rate are projected to be



v = 5.0 and s = 15 percent, respectively.



a. Show that 6 percent growth cannot be achieved under these circumstances. (1 Mark)



b. With the saving rate as given, what ICOR would be required to achieve the 6 percent growth



target? (2 Marks)



c. What happens to the question (a) and (b) if population growth and depreciations in Harrod



Model is ๐’ = ๐œน = ๐ŸŽ. ๐ŸŽ๐Ÿ?

1
Expert's answer
2022-01-31T09:58:52-0500

Estimated growth rate = "5\%"

The ICOR, v = 5.0

The saving rate = "15\%"

a. Show that 4 percent growth cannot be achieved under these circumstances.

"Growth rate = \frac{Saving rate}{ICOR}"

"= \frac{15}{5}
=3.0\%"

Therefore only 3.0 can be achieved under the given conditions and not "6\%"


b.With the saving rate as given, what ICOR would be required to achieve the 6 percent growth target?

"ICOR = \frac{Saving rate}{Growth rate}"

"=\frac{ 15}{6} = 2.5"


c.ฮด stands for the rate of depreciation of the capital stock. For ฮด = 0. 02, the growth rate will be less than "5\%" and the ICOR value above would not be achieved.